OpenStudy (anonymous):

The functions f(x) and g(x) are described using the following equation and table: f(x) = -3(1.02)x x g(x) -1 -5 0 -3 1 -1 2 1 Which statement best compares the y-intercepts of f(x) and g(x)? The y-intercept of f(x) is equal to the y-intercept of g(x). The y-intercept of f(x) is equal to 2 times the y-intercept of g(x). The y-intercept of g(x) is equal to 2 times the y-intercept of f(x). The y-intercept of g(x) is equal to 2 plus the y-intercept of f(x).

OpenStudy (anonymous):

I did but none of them made any sense

undeadknight26 (undeadknight26):

@agent0smith @austinL

OpenStudy (anonymous):

@ganeshie8 @sel95 @radar

OpenStudy (jdoe0001):

I don't see any of the choices fit

OpenStudy (anonymous):

me either thats why its weird

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

@thomaster @TuringTest

OpenStudy (mertsj):

What is the y intercept of the function f(x)?

OpenStudy (the_fizicx99):

This is actually pretty easy, in f(x) = -3(1.02)x -3 is the principle or the y intercept when x is 0. When x is 0: f(0) = -3(1.02)^0 = f(0) = -3(1.02^0) = f(0) = -3(1) = f(0) = -3 Now look at g(x), its also -3 when x is 0. Recall that to find the y intercept you set x for 0 and solve (if there's anything to solve). Now based of this, can you conclude what the answer is @bigoldnastyfish

OpenStudy (radar):

@tHe_FiZiCx99 Where did the 3 come from in f(x) = -3(1.02)x did I miss something? f(x) = -3(1.02)x -3 ??????

OpenStudy (the_fizicx99):

No, -3 is the principle of f(x) = -3(1.02)^x f(x) = P(1+r)^x

OpenStudy (the_fizicx99):

The principle is also the y intercept. Also the start value.

OpenStudy (radar):

I guess I did not believe x was an exponent of 1.02, i was thinking that 1.02(-3) was the coefficient of x. Still stumbling lol

OpenStudy (the_fizicx99):

xD

OpenStudy (anonymous):

@Zale101 @triciaal @goformit100

OpenStudy (anonymous):

@Miracrown

OpenStudy (triciaal):

is this open or closed? I agree the y-intercept is the same (0,-3)

OpenStudy (anonymous):

confuzzled

OpenStudy (triciaal):

anything to the power 0 is 1

OpenStudy (triciaal):

y-intercept is the value of y when x = 0

OpenStudy (triciaal):